OpenStudy (anonymous):

A man bought a TV set that was listed at $160. He was given successive discounts of 20% and 10%. The price he paid was ?

OpenStudy (anonymous):

((($160) * (20% / 100)) * (10% / 100)) = ? Remember 20% is the same thing as twenty 100ths or 0.20, 10% is 0.10

OpenStudy (cwrw238):

a 20% discount means new price is 80% of $160 so multiply 160 by 80/100 then multiply this result by (100-10) = 90% that is 90/100

OpenStudy (anonymous):

cwrw238 wrote it nice steps for you :-)

OpenStudy (cwrw238):

ty agentxS

OpenStudy (anonymous):

you want to say the answer is their product

OpenStudy (anonymous):

the answer should be 115.2

OpenStudy (cwrw238):

(160 * 80 ) / 100 = 160 * 0.8 = $ 128 128 * 0.9 = $ 115.2 note 80/100 = 0.8 and 90/100 = 0.9

OpenStudy (anonymous):

thanks a lot

OpenStudy (cwrw238):

ty - do you understand why we multiplied by 80 % and 90 %?

OpenStudy (anonymous):

yes i get it

OpenStudy (cwrw238):

good - best of luck